### v1

@@ Abstract
Context. Outflows and jets are defining characteristics in protostellar evolution, intimately linked to accretion. Understanding their properties and origins is essential to probing the earliest phases of star formation.
@@
Methods. We obtain spectroscopic observations using JWST MIRI/MRS instrument across m at spatial resolution. We identify emission lines from molecular and ionized species and analyze their spatial morphology using line integrated intensity maps. We derive gas temperature and column density through excitation diagram analysis of H2 rotational lines and compared results with shock models.
Results. The observations reveal extended molecular hydrogen emission tracing the bipolar outflow, with the H2 gas temperatures distributed into warm (550 K) and hot (2500 K) components, likely originating from moderate velocity -type shocks and some UV irradiation. We detect forbidden atomic and ionized emission lines of [Ni ii], [Ar ii], [Ne ii], [Ne iii], [S i], and [Fe ii] showing spatially extended morphology. Double peaked emission profiles were seen in [Ar ii], [Ne iii], and [Fe ii], in the eastern region, suggesting that the high velocity component traces a fast, highly ionized jet. Radial velocity map derived from [Ne ii] emission shows the eastern region to be redshifted and the western region blueshifted, contrary to earlier interpretations.
Conclusions. The analysis of the MIRI/MRS observations reveals the presence of molecular, atomic, and ionized emission lines in this low-mass protostar connected with active outflow signatures. The most striking feature discovered is the presence of a poorly collimated high velocity ionized jet, embedded within a broader wide-angle molecular outflow likely driven by a disk wind. The co-existence of these components supports a stratified outflow structure and suggest L1527 exhibits unique jet-launching characteristics atypical for its early evolutionary stage.
@@ 1 Introduction
The earliest phases of protostar evolution are characterized by accretion and ejection processes, with the forming protostar deeply embedded in its natal envelope. These ejections typically show two connected but distinct observable components: 1) a highly collimated, high-velocity jet (up to several hundred km s-1) whose composition varies with evolutionary stage and may include molecular, atomic, and partially ionized gas, and 2) a slower, wider-angle outflow (km s-1), composed predominantly of molecular and/or atomic gas produced as the jet or a wide-angle wind entrains, shocks, and accelerates the surrounding material (Frank et al., 2014; Bally , 2016; Ray & Ferreira, 2021). The exact mechanism driving these outflows is debated, but the main process involves interaction of accreted matter with the stellar and/or disk magnetic field. Two main models describe the ejection process depending on the distance from the protostar to where the gas is launched. In the X-wind models (Shu et al., 2007; Shang et al., 2007), the launch occurs very close to the protostar where stellar magnetosphere intersects with the protostellar disk. In the Disk-wind models, the winds are launched near to few au’s away from the protostar along the magnetized disk surface (Blandford & Payne , 1982; Shu et al., 1994). In both cases, the poloidal magnetic field plays a critical role in the initial launch and collimation of the outflow (Banerjee & Pudritz , 2006, 2007). It is now understood that outflows are the main mechanism that assist in removal of angular momentum helping accretion to proceed (Ray et al., 2007; Bally , 2016). Recent studies suggest that the slower and wider outflow wind influences the evolution and dispersal of protoplanetary disks (Pascucci et al., 2023). Outflows also play a vital role in limiting the star formation efficiency and shaping the initial mass function (Federrath , 2015; Guszejnov et al., 2022). They can inject energy and momentum into the surrounding medium dispersing a significant amount of the envelope (Dunham et al., 2014; Offner & Arce , 2014; Zhang et al., 2016), thereby ending the infall phase of the protostar.
@@
The outflows hold insights about the associated protostars and can reveal the nature of their origins through their morphologies, kinematics, and physical properties. Outflows can be probed using a wide array of tracers from visual to radio wavelengths. The atomic and ionized components of the outflow have been traced in the visual and near-infrared (NIR) hydrogen recombination lines (Davis et al., 2011; Nisini et al., 2024), and in many forbidden emission transitions such as [O ii], [S ii], [N ii], [Ni ii], [Ca ii], and [Fe ii] (Nisini et al., 2005; Pascucci et al., 2020). The molecular component of outflows have been extensively studied in the infrared and (sub)-millimeter wavelengths using the low-J pure rotational transitions of H2 and CO (Giannini et al., 2002; Lee et al., 2002; Wu et al., 2004; Davis et al., 2010; Mottram et al., 2017; Devaraj et al., 2023). The accelerated warm and hot gas from shocks have been traced in the higher-J transitions of CO (Giannini et al., 2001; Gomez-Ruiz et al., 2012; Kristensen et al., 2017; Karska et al., 2018), and with species such as SiO (Lee , 2020; Towner et al., 2023). More recently in few cases jets have been detected in radio synchrotron emission (Anglada et al., 2018; Feeney-Johansson et al., 2019). However, no tracer is complete, and because of the very high extinction, the deeply embedded regions, particularly the outflows from Class 0/I protostars have to be studied at higher angular resolutions at longer wavelengths. While (sub)-millimeter data from facilities like ALMA have allowed detailed imaging of the cold molecular material (Arce et al., 2013; Hsieh et al., 2023), the mid-infrared (MIR) continuum and spectroscopic imaging of the warmer atomic and molecular material is essential to get a detailed characterization of the processes occurring in the outflows. The James Webb Space Telescope (JWST) (Gardner et al., 2023) with its MIR capabilities can identify and spatially resolve multiple flow components that are hidden at shorter wavelengths, thereby providing a unique window for studying the deeply embedded protostellar phase.
@@
The object in our study is the L1527 IRS (also known as IRAS 04368+2557; hereafter L1527), a low-mass protostar located in the L1527 dark cloud (Beverly , 1962) in the Taurus star-forming region. It is estimated to be at a distance of pc by the most recent studies using Gaia Data Release 2 (DR2) and very long baseline interferometry (VLBI) data (Luhman , 2018; Galli et al., 2019). The protostar has been observed extensively from the near-infrared to centimeter wavelengths. The systemic velocity of the region in the local standard of rest (LSR) frame was estimated to be V km s-1(Ohashi et al., 1997; Tobin et al., 2011). The earliest evidence of outflow in L1527 was identified with (sub)-millimeter observations in CO and HCO+ emissions (Bontemps et al., 1996; Tamura et al., 1996; Zhou et al., 1996; Hogerheijde et al., 1998). Their outflow maps revealed a red shifted and blue shifted emission oriented in east-west direction in the plane-of-sky. A detailed picture of the outflow emission in scattered light, extending 20,000 au, was presented by Tobin et al. (2008, 2010) using Spitzer Space Telescope and Gemini telescope. During Cycle 1 observations, JWST released111(Release ID 2022-055, PI: K. Pontoppidan) https://webbtelescope.org/contents/news-releases/2022/news-2022-055 a high resolution NIR broadband imaging of the outflow, revealing an intricate map of the dust layers and shocked material. Very recently, van’t Hoff et al. (2023) using ALMA data, identified a blue shifted unresolved SiO emission on the western lobe indicating a tentative presence of a jet.
@@
L1527 is often classified as a borderline Class 0/I protostar due its low bolometric temperature (40 K) and the presence of a large and massive envelope (2000 au and 0.9 ) (Jørgensen et al., 2007; Tobin et al., 2013; Ohashi et al., 2023). It is one of the first protostars in its earliest evolutionary stage toward which a Keplerian rotating disk was established (Tobin et al., 2012). The disk has been viewed to be edge-on and its inclination angle was estimated to be (Tobin et al., 2008; Ohashi et al., 2023). High resolution sensitive observations of the disk have shown it to be warped and asymmetric extending up to 100 au (Sheehan et al., 2022; van’t Hoff et al., 2023). Recent analysis of the spectral energy distribution (SED) of L1527 have estimated its bolometric luminosity to be in the range of and a protostar mass of about (Aso et al., 2013; Karska et al., 2018; Ohashi et al., 2023). Using VLA observations at 7 mm, Loinard et al. (2002) directly imaged the disk of L1527 and found an unresolved blob 25 au from the central protostar, suspecting it to be a binary companion. However, many recent high-sensitivity observations with ALMA and VLA have not been able to confirm the binarity (e.g. Nakatani et al., 2020; Sheehan et al., 2022). Although L1527 has been studied at multi-wavelengths presenting an edge-on disk, large envelope mass, and extended outflow cavities, it has not been observed to have any clear collimated high-velocity jet which is typical of Class 0 protostars. This brings into question whether it is steadily accreting in its current state, and if there is at all any definitive evidence of a jet. Therefore, L1527 is an ideal candidate to probe the warm molecular and atomic jet emission using JWST MIR spectroscopy. Additionally, it is thought that the collimation and strength of the jets tends to decline with age of the protostar as the mass accretion rate decreases (e.g. Bontemps et al., 1996; Bally , 2016). Hence, discerning L1527’s outflow properties will be key to understand its evolutionary stage.
@@
In this study, we present new results from JWST MIRI/MRS spectroscopic observations of L1527 covering the inner central region of 500 au. In Section 2, we give a brief overview of the instrument description and observational details. In Section 3, we show the results of the outflow spectra and spatial maps of different emission lines. In Section 4.1, we present physical properties of temperature and column density through excitation diagram analysis of H2 lines with comparison to shock models. In Section 5, we discuss the results and presence of a jet. Finally, in Section 6 we present the summary and concluding remarks.
@@ 2 Observations
L1527 observations with JWST were obtained with the Mid-Infrared Instrument / Medium Resolution Spectrometer (MIRI/MRS; Rieke et al., 2015; Wright et al., 2023; Argyriou et al., 2023). The observations were part of a larger program called “JWST Observations of Young protoStars (JOYS)” (PID1290; van Dishoeck et al., 2025). The JOYS program observes a total of about 20 low to high-mass star-forming cores. The MIRI/MRS instrument does simultaneous spatial and spectral observations between wavelengths of m to m, covering a field of view of up to 7 with a spectral resolving power of to 1500. The instrument has 4 IFU channels with 3 grating sub-band settings (Short, Medium, Long). A full spectrum coverage requires exposure in each of the 3 sub-bands, resulting in 12 spectral segments. MIRI/MRS instrument achieves a PSF FWHM of 0.2′′ at m to 1′′ at m, corresponding to a physical scale of 30 au to 130 au at a distance of 140 pc, thereby providing coverage to the innermost region of L1527.
@@
The science observations were carried out on March 4th, 2023 as part of Cycle 1 JOYS guaranteed time observations. The target position was centered on coordinates 4 +26∘03′0956 (J2000). This pointing is slightly offset from the protostar center (4 +26∘03′0958) on purpose to cover a larger portion of the brighter eastern outflow lobe, but also still observing some part of the western lobe. The data were acquired using FAST1 readout mode with 180 groups of one integration per sub-band. The integration time per sub-band was 1000 s, which resulted in total integration time of 3000 s to obtain the full spectrum. A two-point dither pattern was implemented to cover the extended morphology of the target. Additionally, a dedicated background field was observed with no dithering and having the same target integration time.
@@
The data were reduced from Stage 0 uncalibrated raw files (.uncal) with the JWST pipeline v.1.12.5 (Bushouse et al., 2023), using corresponding Calibration Reference Data System (CRDS) context file jwst1174.pmap. During Stage 1 process, the raw files were processed for detector level corrections using Detector1 Pipeline default settings. This included dark current removal, flagging of bad pixels, non-linearity correction, ramp jump corrections and readout noise removal, resulting in images with count rate per second. During Stage 2 process, the images were calibrated using the Spec2 Pipeline settings. Here, the data was processed for addition of WCS information, flat field correction, background subtraction, flux calibration and fringe correction including removal of residual fringes (Crouzet et al., 2025). An additional bad-pixel removal routine was applied for outlier detection using the Vortex Image Processing (VIP) package (Christiaens et al., 2023). The absolute flux calibration was performed from observations of spectrophotometric standards to convert the detector image to units of MJy sr-1 (Argyriou et al., 2023). Wavelength calibration is applied using MIRI/MRS reference files derived from lab measurements and observations of line emission from astrophysical standards. The wavelength scale is additionally corrected from the observatory reference frame to the barycentric reference frame. Finally, for Stage 3 process, Spec3 Pipeline step was applied which included spectral extraction and cube creation. Totally, 12 data cubes were created corresponding to each of the MIRI/MRS channel and sub-bands. The spectra of all 12 sub-bands were stitched together to allow for a single analysis of the full MIRI/MRS wavelength range.
@@
Since the MIRI spectral cubes are calibrated in the barycentric reference frame, we applied a constant barycentric-to-LSR velocity correction of -10.1 km s-1, appropriate for the coordinates of L1527 and the time of observation. In addition, the velocities were shifted to rest frame by correcting for the systemic velocity of +5.9 km s-1.
@@ 3.1 General MIR imaging and spectral features
In Fig. 1, we present a mid-infrared three-color map of the L1527 region, taken using MIRI broadband filters in m, m and m (PID 1798: J. Tobin & PID 2729: K. Pontoppidan). This map, covering a large field-of-view, allows us to get an improved understanding of the outflow structure in the MIR compared to previous observations. The outflow and the cavity walls are clearly seen as a bright bipolar structure extending to about 6000 au in each direction. Several prominent nebulous filaments consisting of bow shocks, knots and reverse shocks which were unresolved in previous Spitzer MIR images (Tobin et al., 2008), are revealed in this new MIRI map. Some of the bright emission near bow shocks may be associated to UV irradiation. Many cavities are seen near the vicinity of bow shocks where gas and dust is swept away. The outflow is seen to be launched at a wide angle from within the inner 200 au region as previously observed and modeled (Tobin et al., 2008, 2013). Towards the central region, the outflow structure is bright, consisting of the bulk of the emission, whereas towards the wings and rear regions it is fainter. Overall, the outflow appears to be in the plane-of-sky with the eastern flow relatively brighter than the western flow.
@@
The right panel in Fig. 1 shows a zoomed in view of the 500 au central region with brightness scaled to reveal the eastern and western outflow lobes. Both lobes have similar symmetrical morphology, with eastern lobe relatively brighter. The lobes are bisected by a dark lane at the center as expected from the edge-on-disk. The outflow lobes have a bright continuum emission which likely includes scattered light from the inner disk. The position of the forming protostar from millimeter continuum is marked by a star symbol. Two circular apertures in red and blue are shown positioned on the outflow lobes for the MIRI/MRS spectral extraction. The placement of the aperture position was determined such that it is close to the central protostar, but also covering the outflow emission. This would give us a better representation of the properties in the region. The aperture radius was set to , which at m is and at m is . This is twice the size of MIRI FWHM PSF as characterized by Law et al. (2023).
@@
The bottom panel in Fig. 1 shows the full distribution of the MIRI/MRS spectra extracted at the aperture positions. The color-coding of the spectra in red and blue correspond to the eastern and western lobes, respectively. Fig. 2 provides a more detailed view, separating the spectra into four panels corresponding to the four MIRI channels (Ch1, Ch2, Ch3 and Ch4), each observed with three grating settings. In general, the spectra are seen to increase in flux density with wavelength, which is typical for embedded protostars. However, there are several positions with strong absorption features which are characteristics of ices and silicates (Pontoppidan et al., 2008; Öberg et al., 2011; Yang et al., 2022). The ice features comprise of both simple (H2O, CO2, CH4) and complex organic molecules: COM (CH3OH, C2H5OH) (Rocha et al., 2024; van Gelder et al., 2024). The most strongest absorption occur particularly between 5 and m. We also see few positions of sharp dip at around 7.7 and m due to CH4 and CO2 ice (Yang et al., 2022). These solid-state features are indicated by various colored bands in Fig. 2.
@@
At longer wavelengths, the increase in flux density is steep as expected from the warm dust emission. Overall, the eastern and western lobe show a similar spectral energy distribution without any significant difference. Given, the high sensitivity of JWST even with short integration times we detect several spectral lines of molecular, atomic, and ionized emission towards both lobes. Pure rotational H2 (0-0) transitions were detected covering the entire MIRI wavelength range. Forbidden atomic and ionized emission lines of [Ni ii], [Ar ii], [Ne ii], [Ne iii], [S i], and [Fe ii] were also detected. The spectra from the eastern region showed stronger intensities and presence of more ionized emission lines, which are discussed in Section 3.3.

### v2

@@ Abstract
Context. Outflows and jets are defining characteristics in protostellar evolution, intimately linked to accretion. Understanding their properties and origins is essential for probing the earliest phases of star formation.
@@
Methods. We obtained spectroscopic observations using the JWST MIRI/MRS instrument across m at spatial resolution. We identified emission lines from molecular and ionized species and analyzed their spatial morphology using line-integrated intensity maps. We derived gas temperatures and column densities from excitation diagram analysis of H2 rotational lines and compared results with shock models.
Results. The observations reveal extended molecular hydrogen emission tracing the bipolar outflow, with the H2 gas temperatures distributed into warm (550 K) and hot (2500 K) components, likely originating from moderate-velocity -type shocks and some UV irradiation. We detect forbidden atomic and ionized emission lines of [Ni ii], [Ar ii], [Ne ii], [Ne iii], [S i], and [Fe ii] showing spatially extended morphology. Double-peaked emission profiles were seen in [Ar ii], [Ne iii], and [Fe ii] in the eastern region, suggesting that the high-velocity component traces a fast, highly ionized jet. A radial velocity map derived from [Ne ii] emission shows the eastern region to be redshifted and the western region blueshifted, contrary to earlier interpretations.
Conclusions. The analysis of the MIRI/MRS observations reveals molecular, atomic, and ionized emission lines in this low-mass protostar connected with active outflow signatures. The most striking feature discovered is a poorly collimated high-velocity ionized jet, embedded within a broader, wide-angle molecular outflow likely driven by a disk wind. The coexistence of these components supports a stratified outflow structure and suggests that L1527 exhibits jet-launching characteristics atypical of its early evolutionary stage.
@@ 1 Introduction
The earliest phases of protostar evolution are characterized by accretion and ejection processes, with the forming protostar deeply embedded in its natal envelope. These ejections typically show two connected but distinct observable components: 1) a highly collimated, high-velocity (HV) jet (up to several hundred kilometers per second) whose composition varies with evolutionary stage and may include molecular, atomic, and partially ionized gas, and 2) a slower, wider-angle outflow (km s-1), composed predominantly of molecular and/or atomic gas produced as the jet or a wide-angle wind entrains, shocks, and accelerates the surrounding material (Frank et al., 2014; Bally , 2016; Ray & Ferreira, 2021). The exact mechanism driving these outflows is debated, but the main process involves the interaction of accreted matter with the stellar and/or disk magnetic field. Two main models describe the ejection process, depending on the distance from the protostar at which the gas is launched. In the X-wind models (Shu et al., 2007; Shang et al., 2007), the launch occurs very close to the protostar where the stellar magnetosphere intersects with the protostellar disk. In the disk-wind models, the winds are launched from a few au from the protostar along the magnetized disk surface (Blandford & Payne , 1982; Shu et al., 1994). In both cases, the poloidal magnetic field plays a critical role in the initial launch and collimation of the outflow (Banerjee & Pudritz , 2006, 2007). Outflows are now understood to be the primary mechanism in the removal of angular momentum, thereby assisting in accretion (Ray et al., 2007; Bally , 2016). Recent studies suggest that the slower and wider outflow wind influences the evolution and dispersal of protoplanetary disks (Pascucci et al., 2023). Outflows also play a vital role in limiting star formation efficiency and shaping initial mass function (Federrath , 2015; Guszejnov et al., 2022). They can inject energy and momentum into the surrounding medium, dispersing a significant amount of the envelope (Dunham et al., 2014; Offner & Arce , 2014; Zhang et al., 2016) and thereby ending the infall phase of the protostar.
@@
Outflows provide insights into their associated protostars and can reveal the nature of their origins through their morphologies, kinematics, and physical properties. Outflows can be probed using a wide array of tracers from visual to radio wavelengths. The atomic and ionized components of outflows have been traced in visual and near-infrared (NIR) hydrogen recombination lines (Davis et al., 2011; Nisini et al., 2024), and in many forbidden emission transitions such as [O ii], [S ii], [N ii], [Ni ii], [Ca ii], and [Fe ii] (Nisini et al., 2005; Pascucci et al., 2020). The molecular component of outflows have been extensively studied in infrared and (sub)millimeter wavelengths using the low-J pure rotational transitions of H2 and CO (Giannini et al., 2002; Lee et al., 2002; Wu et al., 2004; Davis et al., 2010; Mottram et al., 2017; Devaraj et al., 2023). Accelerated warm and hot gases from shocks have been traced in higher-J transitions of CO (Giannini et al., 2001; Gomez-Ruiz et al., 2012; Kristensen et al., 2017; Karska et al., 2018) and with species such as SiO (Lee , 2020; Towner et al., 2023). More recently, in a few cases, jets have been detected in radio synchrotron emission (Anglada et al., 2018; Feeney-Johansson et al., 2019). However, no tracer is complete, and because of very high extinction, the deeply embedded regions, particularly the outflows from class 0/I protostars, must be studied at higher angular resolutions at longer wavelengths. While (sub)millimeter data from facilities such as ALMA have enabled detailed imaging of cold molecular material (Arce et al., 2013; Hsieh et al., 2023), mid-infrared (MIR) continuum and spectroscopic imaging of warmer atomic and molecular material is essential for a detailed characterization of the outflow processes. The James Webb Space Telescope (JWST) (Gardner et al., 2023) with its MIR capabilities can identify and spatially resolve multiple flow components hidden at shorter wavelengths, providing a unique window into the deeply embedded protostellar phase.
@@
The object of our study is L1527 IRS (also known as IRAS 04368+2557; hereafter L1527), a low-mass protostar located in the L1527 dark cloud (Beverly , 1962) in the Taurus star-forming region. It is estimated to lie at a distance of pc by the most recent studies using Gaia data release 2 (DR2) and very long baseline interferometry (VLBI) data (Luhman , 2018; Galli et al., 2019). The protostar has been observed extensively from the NIR to centimeter wavelengths. The systemic velocity of the region in the Local-Standard-of-Rest (LSR) frame was estimated to be V km s-1(Ohashi et al., 1997; Tobin et al., 2011). The earliest evidence of outflow in L1527 was identified with (sub)millimeter observations in CO and HCO+ emissions (Bontemps et al., 1996; Tamura et al., 1996; Zhou et al., 1996; Hogerheijde et al., 1998). Their outflow maps revealed a redshifted and blueshifted emission oriented in the east-west direction in the plane of sky. A detailed picture of the outflow emission in scattered light, extending 20,000 au, was presented by Tobin et al. (2008, 2010) using the Spitzer Space Telescope and Gemini telescope. During Cycle 1 observations, JWST released111(Release ID 2022-055, PI: K. Pontoppidan) https://webbtelescope.org/contents/news-releases/2022/news-2022-055 high-resolution NIR broadband imaging of the outflow, revealing an intricate map of its dust layers and shocked material. Using ALMA data, van’t Hoff et al. (2023) very recently identified a blueshifted unresolved SiO emission on the western lobe, indicating a tentative jet.
@@
L1527 is often classified as a borderline class 0/I protostar due its low bolometric temperature (40 K) as well as its large and massive envelope (2000 au and 0.9 ) (Jørgensen et al., 2007; Tobin et al., 2013; Ohashi et al., 2023). It is one of the first protostars in its earliest evolutionary stage toward which a Keplerian rotating disk was established (Tobin et al., 2012). The disk has been viewed to be edge-on, and its inclination angle was estimated to be (Tobin et al., 2008; Ohashi et al., 2023). High-resolution sensitive observations of the disk have shown it to be warped and asymmetric, extending up to 100 au (Sheehan et al., 2022; van’t Hoff et al., 2023). A recent analysis of L1527’s spectral energy distribution (SED) has estimated its bolometric luminosity to be in the range of and a protostar mass of about (Aso et al., 2013; Karska et al., 2018; Ohashi et al., 2023). Using VLA observations at 7 mm, Loinard et al. (2002) directly imaged the disk of L1527 and found an unresolved blob 25 au from the central protostar, suspecting it to be a binary companion. However, many recent high-sensitivity observations with ALMA and VLA have not been able to confirm binarity (e.g., Nakatani et al., 2020; Sheehan et al., 2022). Although L1527 has been studied at several wavelengths – revealing an edge-on disk, large envelope mass, and extended outflow cavities – it has not been observed to have any clear collimated HV jet typical of class 0 protostars. This raises the question of whether it is steadily accreting in its current state and if definitive evidence of a jet exists at all. Therefore, L1527 is an ideal candidate for probing warm molecular and atomic jet emission using JWST MIR spectroscopy. Additionally, it is thought that the collimation and strength of jets decline with protostar age as the mass accretion rate decreases (e.g., Bontemps et al., 1996; Bally , 2016). Hence, discerning L1527’s outflow properties will be key to understanding its evolutionary stage.
@@
In this study, we present new results from JWST MIRI/MRS spectroscopic observations of L1527 covering the inner central region of 500 au. In Section 2, we briefly describe the instrument and provide observational details. In Section 3, we present the results of outflow spectra and spatial maps of different emission lines. In Section 4.1, we present the physical properties (temperature and column density) derived from excitation diagram analysis of H2 lines and their comparison with shock models. In Section 5, we discuss the results and presence of a jet. Finally, in Section 6 we present the summary and concluding remarks.
@@ 2 Observations
L1527 observations with JWST were obtained with the Mid-Infrared Instrument/Medium Resolution Spectrometer (MIRI/MRS; Rieke et al., 2015; Wright et al., 2023; Argyriou et al., 2023). The observations were part of a larger program called “JWST Observations of Young protoStars (JOYS)” (PID1290; van Dishoeck et al., 2025). The JOYS program observes a total of about 20 low- to high-mass star-forming cores. The MIRI/MRS instrument does simultaneous spatial and spectral observations between wavelengths of m to m, covering a field of view of up to 7 with a spectral resolving power of to 1500. The instrument has four IFU channels with three grating sub-band settings (short, medium, and long). A full spectrum coverage requires exposure in each of the three sub-bands, resulting in 12 spectral segments. MIRI/MRS achieves a point spread function (PSF) full width at half maximum (FWHM) of 0.2′′ at m to 1′′ at m, corresponding to a physical scale of 30 au to 130 au at a distance of 140 pc, thereby providing coverage to the innermost region of L1527.
@@
Science observations were carried out on March 4, 2023, as part of Cycle 1 JOYS guaranteed time observations. The target position was centered on coordinates 4 +26∘03′0956 (J2000). This pointing was slightly offset from the protostar center (4 +26∘03′0958) to cover more of the brighter eastern outflow lobe while still observing part of the western lobe. The data were acquired using FAST1 readout mode with 180 groups of one integration per sub-band. The integration time per sub-band was 1000 s, which resulted in a total integration time of 3000 s to obtain the full spectrum. A two-point dither pattern was implemented to cover the extended morphology of the target. Additionally, a dedicated background field was observed with no dithering and with the same target integration time.
@@
The data were reduced from Stage 0 uncalibrated raw files (.uncal) with the JWST pipeline v.1.12.5 (Bushouse et al., 2023), using corresponding the Calibration Reference Data System (CRDS) context file jwst1174.pmap. During Stage 1 processing, the raw files were processed for detector-level corrections using Detector1 Pipeline default settings. This included dark current removal, bad pixel flagging, nonlinearity correction, ramp jump corrections, and readout noise removal, resulting in images with count rates per second. During Stage 2 processing, the images were calibrated using the Spec2 Pipeline settings. Here, the data underwent WCS addition, flat-field correction, background subtraction, flux calibration, and fringe correction including residual fringe removal (Crouzet et al., 2025). An additional bad-pixel removal routine was applied for outlier detection using the Vortex Image Processing (VIP) package (Christiaens et al., 2023). Absolute flux calibration was performed from observations of spectrophotometric standards to convert the detector image to units of megaJanskys per steradian (Argyriou et al., 2023). Wavelength calibration was applied using MIRI/MRS reference files derived from lab measurements and observations of line emission from astrophysical standards. The wavelength scale was additionally corrected from the observatory reference frame to the barycentric reference frame. Finally, for Stage 3 processing, the Spec3 Pipeline step was applied, which included spectral extraction and cube creation. In total, 12 data cubes were created, corresponding to each of the MIRI/MRS channel and sub-bands. The spectra of all 12 sub-bands were stitched together to enable a single analysis of the full MIRI/MRS wavelength range.
@@
Since the MIRI spectral cubes were calibrated in the barycentric reference frame, we applied a constant barycentric-to-LSR velocity correction of -10.1 km s-1, appropriate for L1527 coordinates and the observation time. In addition, the velocities were shifted to the rest frame by correcting for the systemic velocity of +5.9 km s-1.
@@ 3.1 General MIR imaging and spectral features
In Fig. 1, we present a MIR three-color map of the L1527 region, taken using MIRI broadband filters in m, m, and m (PID 1798: J. Tobin and PID 2729: K. Pontoppidan). This map, covering a large field of view, provides improved understanding of the outflow structure in the MIR compared to previous observations. The outflow and cavity walls are clearly seen as a bright bipolar structure extending to about 6000 au in each direction. Several prominent nebulous filaments, consisting of bow shocks, knots, and reverse shocks that were unresolved in previous Spitzer MIR images (Tobin et al., 2008), are revealed in this new MIRI map. Some of the bright emission near the bow shocks may be associated with UV irradiation. Many cavities appear near the vicinity of the bow shocks where gas and dust are swept away. The outflow is launched at a wide angle from within the inner 200 au region as previously observed and modeled (Tobin et al., 2008, 2013). Toward the central region, the outflow structure is bright and contains the bulk of the emission, whereas it is fainter toward the wings and rear regions. Overall, the outflow appears to be in the plane of sky with the eastern flow relatively brighter than the western flow.
@@
The right panel in Fig. 1 shows a zoomed in view of the 500 au central region with brightness scaled to reveal the eastern and western outflow lobes. Both lobes exhibit similar symmetrical morphology, with the eastern lobe being relatively brighter. The lobes are bisected by a dark lane at the center, as expected from the edge-on disk. The outflow lobes show bright continuum emission, which likely includes scattered light from the inner disk. The position of the forming protostar from millimeter continuum is marked by a star. Two circular apertures in red and blue are shown positioned on the outflow lobes for MIRI/MRS spectral extraction. These apertures were placed close to the central protostar while covering outflow emission, better representing the region’s properties. The aperture radius was set to , which gives at m and at m. This is twice the MIRI PSF FWHM size as characterized by Law et al. (2023).
@@
The bottom panel in Fig. 1 shows the full distribution of the MIRI/MRS spectra extracted at the aperture positions. The color-coding of the spectra in red and blue correspond to the eastern and western lobes, respectively. Fig. 2 provides a more detailed view, separating the spectra into four panels corresponding to the four MIRI channels (Ch1, Ch2, Ch3, and Ch4), each observed with three grating settings. In general, the spectra increase in flux density with wavelength, which is typical for embedded protostars. However, several positions show strong absorption features characteristic of ices and silicates (Pontoppidan et al., 2008; Öberg et al., 2011; Yang et al., 2022). The ice features comprise both simple (H2O, CO2, and CH4) and complex organic molecules (COM (CH3OH, C2H5OH); (Rocha et al., 2024; van Gelder et al., 2024)). Strongest absorption appears particularly between 5 and m. We also see sharp dips at a few positions around 7.7 and m due to CH4 and CO2 ice (Yang et al., 2022). These solid-state features are indicated by various colored bands in Fig. 2.
@@
At longer wavelengths, the increase in flux density is steep, as expected from the warm dust emission. Overall, the eastern and western lobes show a similar SED without any significant difference. Given the high sensitivity of JWST even with short integration times, we detect several spectral lines of molecular, atomic, and ionized emission toward both lobes. Pure rotational H2 (0-0) transitions were detected covering the entire MIRI wavelength range. Forbidden atomic and ionized emission lines of [Ni ii], [Ar ii], [Ne ii], [Ne iii], [S i], and [Fe ii] were also detected. The spectra from the eastern region showed stronger intensities as well as more ionized emission lines, which are discussed in Section 3.3.